What is the potential difference VB - VA when I =0.50 A in the circuit segmentshown below?
Answer #1
Going from B to A you...
   drop potential across the first resistorof   V = IR = 0.5 * 16 = 8volts
   drop potential across the battery (go from + sideto - side) of    15 V
drop potential across second resistorof    IR = 0.50 * 10 =   5volts
If you think of this like going downhill, you started atthe top at point B and went downhill
     8   +  15   + 5 =   28 Volts
So B is higher than A by 28 volts   and thedifference of VB - VA is + 28